The admission fee at an amusement park is 1.5 dollars for children 4 dollars for adults. On a certain day 315 people entered the park, and the admission fees collected 810 dollars. How many children and how many adults were admitted?
Answer by checkley77(12569)

A+C=315 OR A=315-C
4A+1.5C=810
4(315-C)+1.5C=810
1,260-4C+1.5C=810
-2.5C=810-1,260
-2.5C=-450
C=-450/-2.5
C=180 CHILDREN TICKETS WERE SOLD.
315-180=135 ADULT TICKETS WERE SOLD.
PROOF:
4*135+1.5*180=810
540+270=810
810=810